# Spectral instability of horizonless compact objects within astrophysical environments: The “exotic” elephant and the flea

### v1

@@ I Introduction
Gravitational-wave (GW) astrophysics has entered a transformative era since the landmark detection by LIGO in 2015 [1], followed by the most recent and loudest event in 2025 [2], thus confirming a century-old prediction of Einstein’s General Relativity (GR). Current state-of-the-art efforts involve a global network of detectors, including LIGO, Virgo, and KAGRA, operating with ever-increasing sensitivity to capture faint ripples in spacetime caused by cataclysmic astrophysical events like black hole (BH) and neutron star mergers [3]. These observations have opened a new window into the universe, enabling precision tests of GR in extreme regimes and offering insights into the population, formation, and evolution of compact objects [4, 5]. Meanwhile, next-generation, observatories, such as the Laser Interferometer Space Antenna [6], the Einstein Telescope [7, 8, 9], and Cosmic Explorer [10, 11] among others, promise to expand the observable sensitivity and frequency range onto uncharted regions of the Cosmos [12, 13, 14, 15, 16, 17, 18, 19, 20]. Thus, they will provide access deeper into the early universe [21, 22, 23, 24], uncover beyond the Standard Model physics, such as dark matter [25, 26, 27, 28], perceive the role of astrophysical environments around BHs [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57] and unearth dark exotic compact objects (ECOs) that are currently theorized [58, 59, 60, 61, 62, 63, 64, 65, 66, 67].
@@
Recently, the mathematical tools of non-modal analysis, i.e., the study of (non-)selfadjoint operators without the use of eigenvalues, appeared in the field of BH perturbation theory [152]. It turns out that the non-selfadjoint operator that describes wave propagation in fixed curved spacetimes and, in particular, the QNMs of compact objects are exponentially sensitive to tiny fluctuations to the effective potential; a phenomenon known as spectral instability [153, 154, 155, 156, 157, 152, 158, 159, 160, 161]. The fragility of the BH spectra can be exposed through tiny alterations of the boundary conditions [76, 162, 163, 164], by the presence of astrophysical configurations [159, 165, 166, 167, 168, 55] or by near-horizon deviations from the BH paradigm [59, 60, 61, 169]. Nevertheless, the prompt ringdown signal is either much less affected by this apparent QNM instability [168, 170, 171, 172], or other quantities indirectly connected to the ringdown appear to be spectrally-stable [173, 174, 175, 176, 177, 160].
@@ III.1 Hyperboloidal approach to ECO QNMs
The final matrix, $\mathbf{L}_{T}$, has dimensions of $2(N+1)\times 2(N+1)$. The transition conditions are imposed by modifying two rows in $\mathbf{L}_{T}$, which results in a single row with at least one non-zero value in the $\mathbf{BC}_{1,2}$ matrices, and in $\mathbf{B}$ which is an identity matrix, besides the two corresponding rows which must be zero. These transition conditions impose the continuity of the function and its derivative at the bump location, and the Dirichlet condition imposes the ECO condition, such that solving this generalized eigenvalue problem, we find the QNMs of the system. The data generated and presented in the rest of this work have been obtain with $N=300$ collocation points to ensure numerical convergence.
@@ IV The role of the ECO compactness on QNMs
Interpreting the absence of a spectral instability with respect to the compactness of the ECO can be realized in two ways: (i) the ECO acts as a typical object where its dynamics continuously change when we vary the available parameters. Note though that the surface position is added by hand and does not backreact to the spacetime metric, which is otherwise Schwarzschild besides the boundary condition at the ECO surface. (ii) In a more theoretical and counter-intuitive level of understanding spectral instabilities, and their connection with the pseudospectrum shown in Ref. [183], we are utilizing an object that has already undergone a drastic change of features with respect to its exterior BH counterpart. The change of boundary condition at the ECO surface can be considered equivalent to adding a very large fluctuation to the BH effective potential, in the form of either high-frequency sinusoids or random perturbations. Therefore, the ECO modes can be considered as a dramatically destabilized set of their BH QNM counterpart, where a very large perturbation has been introduced to the effective potential, in the form of a complete alteration of the near-horizon boundary condition.
@@ V.1 Destabilization of ECO QNMs
Figures 3 and 4 show collectively the real and imaginary part of the fundamental mode and first overtone for three horizonless ECO configurations with compactness $\mathcal{E}/r_{h}=10^{-3},\,10^{-2}$ and $10^{-1}$, where the bump location $a_{0}/r_{h}$ serves as a free parameter. Firstly, we observe that the fundamental mode of ECOs with compactness $\mathcal{E}/r_{h}=10^{-3},\,10^{-2}$ remains practically unchanged, regardless of how far (even beyond an astrophysical sense) we place the bump. Thus, compact ($\mathcal{E}/r_{h}=10^{-2}$) and ultra-compact ($\mathcal{E}/r_{h}=10^{-3}$) exotic horizonless objects have spectrally-stable fundamental QNMs, for a sufficiently large range of $a_{0}/r_{h}$. On the other hand, when the compactness decreases, i.e., the ECO surface is further away from the Schwarzschild radius ($\mathcal{E}/r_{h}=10^{-1}$), then a migration of the fundamental mode occurs in the complex plane, since beyond $a_{0}\sim 25r_{h}$ both the real and imaginary parts of the fundamental mode change by $\mathcal{O}(10^{-1})$ and $\mathcal{O}(10^{-2})$, respectively. This is a clear occurrence of spectral instability since the lowest-lying mode migrates disproportionately in the complex plane if one considers that the Gaussian bump has an amplitude or order $\mathcal{O}(10^{-6})$ [^3] See though the discussion in Ref. [166].. The existence of these new “perturbed” modes, that may destabilize the spectrum, is due to the creation of a secondary trapping cavity in the exterior of the light ring, between it and the environmental bump. We dub these QNMs as exterior trapped modes.
@@ V.2 The nature of the overtaking instability
As Fig. 6 designates that for a fixed compactness, the ECO QNMs are spectrally-destabilized, through an overtaking instability, beyond some bump location, so Fig. 7 shows the phase space ($\mathcal{E}$, $a_{0}$) where spectral migration and overtaking instability from interior to exterior modes occurs. For example, when $\mathcal{E}/r_{h}=0.3$ the fundamental mode is overtaken for a bump location at $a_{0}\gtrsim 15r_{h}$. On the other hand, to destabilize the first overtone of the ECO, with a bump at $a_{0}\gtrsim 15r_{h}$, a compactness $\mathcal{E}/r_{h}=0.025$ is only needed. Similarly, when $a_{0}\gtrsim 60r_{h}$ then a smaller compactness (larger $r_{s}$), i.e., $\mathcal{E}/r_{h}=0.04$ is needed for the fundamental mode to be overtaken and $\mathcal{E}/r_{h}=0.004$ is needed for the same result on the first overtone. Hence, the more compact the ECO is (smaller surface radius $r=r_{s}$), the further we need to place the environmental bump from the light ring in order to destabilize the spectrum. Figures 6 and 7 uncover the fact that the QNMs of ultra-compact ECOs present a “strong” spectral stability, until they are eventually destabilized with environmental effects, since the bump needs to be located sufficiently far from the light ring of the object. In a similar sense, QNMs from less-compact ECOs (larger surface radius $r=r_{s}$) can be destabilized “easier” by placing a bump in a location closer to the light ring, with respect to ultra-compact configurations. Loosely-compact ECOs with reflective surfaces and a light ring are more prone to spectral instability due to environmental bumps than ultra-compact ones.
@@
The abrupt metamorphosis from a Schwarzschild BH to a horizonless object that we imposed through the reflective boundary condition at the surface, combined with the successive overtaking spectral instability due to environmental effects in non-compact ECOs demonstrate that these objects exhibit both a strong destabilization of their interior, as well as the exterior modes that repetitively takes place, and acquire smaller (in absolute value) imaginary parts with respect to the $a_{0}=0$ unperturbed QNM. This leads to a series of new “perturbed” fundamental QNMs as $a_{0}/r_{h}$ is increased. Thus, ECOs present spectral instabilities of both their overtones and fundamental mode, depending on the compactness $\mathcal{E}/r_{h}$. It is important to note that ultra-compact configurations need radial bump locations that may possibly be too large and, therefore, astrophysically irrelevant. Thus, one could say that the fundamental mode of ultra-compact ECOs cannot be destabilized, in astrophysical scenarios.
@@ V.3 Transition from spectral to modal instability?
Finally, we note that in this work, we have taken into account idealized, spherically-symmetric ECOs with a perfectly reflective surface. Nevertheless, nature abhors exactly spherically-symmetric configurations. BHs and compact objects in our Universe are rotating, thus are stationary and axially symmetric. These BHs solutions, like the Kerr spacetime, or rotating ECOs [130, 133, 137, 62, 218, 203, 126] have more intricate spectra and only a preliminary analysis of the pseudospectrum of Kerr BHs has been achieved, for scalar field perturbations [191]. Therefore, possible new phenomena that may induce spectral-to-modal instabilities cannot be eliminated till rotation is included to the study performed here. The same holds for the form of the perturbation induced through environmental effects. Here, we have used smooth Gaussian bumps, that possess regularity proportional to their sample paths in terms of their derivatives. Perhaps, another Gaussian bump with lower regularity and smoothness might trigger spectral instabilities that lead to the destabilization of the ECO spacetime itself.
@@ Acknowledgements.
K.D. acknowledges financial support provided by FCT – Fundação para a Ciência e a Tecnologia, I.P., under the Scientific Employment Stimulus – Individual Call – Grant No. 2023.07417.CEECIND/CP2830/CT0008. M.M.C. and C.F.B.M. acknowledge Fundação Amazônia de Amparo a Estudos e Pesquisas (FAPESPA), Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) – Finance Code 001, from Brazil, for partial financial support. R.P.M. acknowledges support from the Villum Investigator program supported by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. This project has received funding from the European Union’s Horizon MSCA-2022 research and innovation programme “Einstein Waves” under grant agreement No. 101131233.

### v2

@@ I Introduction
Gravitational-wave (GW) astrophysics has entered a transformative era since the landmark detection by LIGO in 2015 [1], followed by the most recent and loudest event in 2025 [2], thus confirming a century-old prediction of Einstein’s General Relativity (GR). Current state-of-the-art efforts involve a global network of detectors, including LIGO, Virgo, and KAGRA, operating with ever-increasing sensitivity to capture faint ripples in spacetime caused by cataclysmic astrophysical events like black hole (BH) and neutron star mergers [3]. These observations have opened a new window into the universe, enabling precision tests of GR in extreme regimes and offering insights into the population, formation, and evolution of compact objects [4, 5]. Meanwhile, next-generation, observatories, such as the Laser Interferometer Space Antenna [6], the Einstein Telescope [7, 8, 9], and Cosmic Explorer [10, 11] among others, promise to expand the observable sensitivity and frequency range onto uncharted regions of the Cosmos [12, 13, 14, 15, 16, 17, 18, 19, 20]. Thus, they will provide access deeper into the early universe [21, 22, 23, 24], uncover beyond the Standard Model physics, such as dark matter [25, 26, 27, 28], perceive the role of astrophysical environments around BHs [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57] and unearth dark exotic compact objects (ECOs) that are currently theorized [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68]. The pay-off for facing these outstanding issues, regarding the existence of ECOs, is to be able to quantify the statement that BHs exist in nature.
@@
Recently, the mathematical tools of non-modal analysis, i.e., the study of (non-)selfadjoint operators without the use of eigenvalues, appeared in the field of BH perturbation theory [153]. It turns out that the non-selfadjoint operator that describes wave propagation in fixed curved spacetimes and, in particular, the QNMs of compact objects are exponentially sensitive to tiny fluctuations to the effective potential; a phenomenon known as spectral instability [154, 155, 156, 157, 158, 153, 159, 160, 161, 162]. The fragility of the BH spectra can be exposed through tiny alterations of the boundary conditions [77, 163, 164, 165], by the presence of astrophysical configurations [160, 166, 167, 168, 169, 55] or by near-horizon deviations from the BH paradigm [60, 61, 62, 170]. Nevertheless, the prompt ringdown signal is either much less affected by this apparent QNM instability [169, 171, 172, 173], or other quantities indirectly connected to the ringdown appear to be spectrally-stable [174, 175, 176, 177, 178, 161, 179].
@@ II.2 Scalar perturbations and quasinormal modes
In principle, a horizonless compact object with pure reflectivity surrounded by a matter halo (that introduces the “flea”), could be constructed with the Einstein cluster formalism. For example, in the supplemental material of Ref. [160], where a procedure is described that embeds a non-thin matter halo to any spherically-symmetric spacetime is presented. What would change in our current object is its reflectivity at its surface. To surpass this issue we could utilize a spherically-symmetric ansatz, such as that of a gravastar or a thin matter shell [61], which have well-defined boundary conditions at their surfaces. In fact, similar constructions have been found, for example, in [115] where an exotic compact object is constructed to have a matter shell around it, in order to understand how echoes of the surface of the ECO are affected by the surrounding astrophysical environment. Also, since it is now widely established that globular clusters host populations of white dwarf, neutron star, and BH binaries, dynamical processes enabled by stellar densities millions of times larger than typical galactic environments, facilitate interactions involving these stellar remnants that give rise to an array of astrophysical phenomena [206]. Nevertheless, for a full understanding of astrophysical environments and their interaction with ultra-compact objects, the development of a solid theoretical framework should be developed. Since a proper modeling of the coalescence, ringing and data analysis of such objects is mostly lacking, we use the simplified analysis above to gain an initial insight.
@@ III.1 Hyperboloidal approach to ECO QNMs
Figure 2: Top row: Convergence tests of the fundamental $\ell=2$ mode ($n=0$, top left) and first overtone ($n=1$, top right) for an ECO with $\mathcal{E}/r_{h}=10^{-3}$ and a bump at varying positions $a_{0}$ and amplitude $\epsilon=10^{-6}$. Due to the absence of of an exact spectrum, we have used the $N=300$-grid-points value of QNMs as our baseline and compared its convergence by varying NN. We find that the absolute difference of the imaginary parts between the $N=300$ and, say, $N=250$ is of order $\mathcal{O}(10^{-3}-10^{-4})$. The same order of magnitudes are found for the real part of the QNMs compared. Bottom row: Same as above but for an ECO with $\mathcal{E}/r_{h}=10^{-2}$. We find that the absolute difference of the imaginary parts between the $N=300$ and, say, $N=250$ is of order $\mathcal{O}(10^{-8}-10^{-10})$ for the fundamental mode (bottom left) and $\mathcal{O}(10^{-4}-10^{-8})$. The same order of magnitudes are found for the real part of the QNMs compared. The exponential convergence in all plots confirms the expected behavior from the employed spectral scheme.
@@
The final matrix, $\mathbf{L}_{T}$, has dimensions of $2(N+1)\times 2(N+1)$. The transition conditions are imposed by modifying two rows in $\mathbf{L}_{T}$, which results in a single row with at least one non-zero value in the $\mathbf{BC}_{1,2}$ matrices, and in $\mathbf{B}$ which is an identity matrix, besides the two corresponding rows which must be zero. These transition conditions impose the continuity of the function and its derivative at the bump location, and the Dirichlet condition imposes the ECO condition, such that solving this generalized eigenvalue problem, we find the QNMs of the system. The data generated and presented in the rest of this work have been obtain with $N=300$ collocation points to ensure numerical convergence, as shown in Fig. 2. In fact, it is noteworthy to mention that the convergence increases rapidly when the ECO is less compact.
@@ IV The role of the ECO compactness on QNMs
Interpreting the absence of a spectral instability with respect to the compactness of the ECO can be realized in two ways: (i) the ECO acts as a typical object where its dynamics continuously change when we vary the available parameters. Note though that the surface position is added by hand and does not backreact to the spacetime metric, which is otherwise Schwarzschild besides the boundary condition at the ECO surface. (ii) In a more theoretical and counter-intuitive level of understanding spectral instabilities, and their connection to the pseudospectrum, shown in Ref. [185], we are utilizing an object that has already undergone a drastic change of features with respect to its exterior BH counterpart. In other words, contrary to the QNM spectral instability paradigm, where small perturbations in the system lead to a large deviation in the spectrum, the ECO modes can be considered as a dramatically destabilized spectral set with respect to their BH QNM counterpart, because a very large modification has been introduced to the system, in the form of a complete alteration of the near-horizon boundary condition. However, once the new QNMs arise, the spectra is stable with respect to small deviation of the parameter controlling the location of the boundary condition.
@@ V.1 Destabilization of ECO QNMs
Figures 4 and 5 show collectively the real and imaginary part of the fundamental mode and first overtone for three horizonless ECO configurations with compactness $\mathcal{E}/r_{h}=10^{-3},\,10^{-2}$ and $10^{-1}$, where the bump location $a_{0}/r_{h}$ serves as a free parameter. Firstly, we observe that the fundamental mode of ECOs with compactness $\mathcal{E}/r_{h}=10^{-3},\,10^{-2}$ remains practically unchanged, regardless of how far (even beyond an astrophysical sense) we place the bump. Thus, compact ($\mathcal{E}/r_{h}=10^{-2}$) and ultra-compact ($\mathcal{E}/r_{h}=10^{-3}$) exotic horizonless objects have spectrally-stable fundamental QNMs, for a sufficiently large range of $a_{0}/r_{h}$. On the other hand, when the compactness decreases, i.e., the ECO surface is further away from the Schwarzschild radius ($\mathcal{E}/r_{h}=10^{-1}$), then a migration of the fundamental mode occurs in the complex plane, since beyond $a_{0}\sim 25r_{h}$ both the real and imaginary parts of the fundamental mode change by $\mathcal{O}(10^{-1})$ and $\mathcal{O}(10^{-2})$, respectively. This is a clear occurrence of spectral instability since the lowest-lying mode migrates disproportionately in the complex plane if one considers that the Gaussian bump has an amplitude or order $\mathcal{O}(10^{-6})$ [^3] See though the discussion in Ref. [167].. The existence of these new “perturbed” modes, that may destabilize the spectrum, is due to the creation of a secondary trapping cavity in the exterior of the light ring, between it and the environmental bump. We dub these QNMs as exterior trapped modes.
@@ V.2 The nature of the overtaking instability
As Fig. 7 designates that for a fixed compactness, the ECO QNMs are spectrally-destabilized, through an overtaking instability, beyond some bump location, so Fig. 8 shows the phase space ($\mathcal{E}$, $a_{0}$) where spectral migration and overtaking instability from interior to exterior modes occurs. For example, when $\mathcal{E}/r_{h}=0.3$ the fundamental mode is overtaken for a bump location at $a_{0}\gtrsim 15r_{h}$. On the other hand, to destabilize the first overtone of the ECO, with a bump at $a_{0}\gtrsim 15r_{h}$, a compactness $\mathcal{E}/r_{h}=0.025$ is only needed. Similarly, when $a_{0}\gtrsim 60r_{h}$ then a smaller compactness (larger $r_{s}$), i.e., $\mathcal{E}/r_{h}=0.04$ is needed for the fundamental mode to be overtaken and $\mathcal{E}/r_{h}=0.004$ is needed for the same result on the first overtone. Hence, the more compact the ECO is (smaller surface radius $r=r_{s}$), the further we need to place the environmental bump from the light ring in order to destabilize the spectrum. Figures 7 and 8 uncover the fact that the QNMs of ultra-compact ECOs present a “strong” spectral stability, until they are eventually destabilized with environmental effects, since the bump needs to be located sufficiently far from the light ring of the object. In a similar sense, QNMs from less-compact ECOs (larger surface radius $r=r_{s}$) can be destabilized “easier” by placing a bump in a location closer to the light ring, with respect to ultra-compact configurations. Loosely-compact ECOs with reflective surfaces and a light ring are more prone to spectral instability due to environmental bumps than ultra-compact ones.
@@
The abrupt metamorphosis from a Schwarzschild BH to a horizonless object that we imposed through the reflective boundary condition at the surface, combined with the successive overtaking spectral instability due to environmental effects in non-compact ECOs demonstrate that these objects exhibit both a strong destabilization of their interior, as well as the exterior modes that repetitively takes place, and acquire smaller (in absolute value) imaginary parts with respect to the $a_{0}=0$ unperturbed QNM. This leads to a series of new “perturbed” fundamental QNMs as $a_{0}/r_{h}$ is increased. Thus, ECOs present spectral instabilities of both their overtones and fundamental mode, depending on the compactness $\mathcal{E}/r_{h}$. It is important to note that ultra-compact configurations need radial bump locations that may possibly be too large and, therefore, astrophysically irrelevant. Thus, one could say that the fundamental mode of ultra-compact ECOs cannot be destabilized, in astrophysical scenarios. On the other hand, loosely-compact, purely-reflective, ECOs, even though slightly more astrophysically relevant, present strong destabilization of the fundamental QNM. Such ECOs are still unrealistic due to their purely reflective surface.
From a thought experiment, where the purely reflective surface transformed slowly to a surface that is partially reflective (and thus also partially absorbing), we may extrapolate that the bump can be placed even further in order to give rise to exterior longer-lived modes that would eventually destabilize the fundamental mode. This is in agreement with the fact that the Schwarzschild BH QNMs can be destabilized completely without the need of near-horizon alterations (see Ref. [160]). Thus, the bump is the main source of the spectral instability of QNMs and not the percentage of reflectivity of the ECO surface. Of course, more exploration is needed, where the reflectivity is also altered.
We emphasize that, from a physical standpoint, rotation could in principle destabilize the ECO, potentially invalidating the physical relevance of our modes. However, Ref. [131] demonstrated that even a small amount of partial absorption is sufficient to stabilize the system. Such absorption can be interpreted as arising from physical mechanisms, for instance effective viscosity (see [219]).
@@ V.3 Transition from spectral to modal instability?
Finally, we note that in this work, we have taken into account idealized, spherically-symmetric ECOs with perfectly reflective surfaces and an astrophysical matter bump. Nevertheless, nature abhors exactly spherically-symmetric configurations. BHs and compact objects in our Universe are rotating, thus are stationary and axially symmetric. These BHs solutions, like the Kerr spacetime, or rotating ECOs [131, 134, 138, 63, 221, 205, 127] have more intricate spectra and only a preliminary analysis of the pseudospectrum of Kerr BHs has been achieved, for scalar field perturbations [193]. Therefore, possible new phenomena that may induce spectral-to-modal instabilities cannot be eliminated till rotation is included to the study performed here. The same holds for the form of the perturbation induced through environmental effects.
Here, we have used smooth Gaussian bumps, that possess regularity proportional to their sample paths in terms of their derivatives. Perhaps, another Gaussian bump with lower regularity and smoothness might trigger spectral instabilities that lead to the destabilization of the ECO spacetime itself. Nevertheless, for all tests we have performed yet, i.e., increasing the amplitude of the bump $\epsilon$ (up to $\mathcal{O}(10^{-3})$), the distance of the bump from the light ring $a_{0}/r_{h}$ (up to almost unrealistic values), and the index $\ell$, that contributes to more modes migrating to the real axis.
@@ Acknowledgements.
K.D. acknowledges financial support provided by FCT – Fundação para a Ciência e a Tecnologia, I.P., under the Scientific Employment Stimulus – Individual Call – Grant No. 2023.07417.CEECIND/CP2830/CT0008. K.D. acknowledges financial support provided by FCT–Fundação para a Ciência e a Tecnologia, I.P., under the Scientific Employment Stimulus – Individual Call – Grant No. 2023.07417.CEECIND/CP2830/CT0008. K.D. would also like to thank the Fundação para a Ciência e Tecnologia (FCT), Portugal, for the financial support to the Center for Astrophysics and Gravitation (CENTRA/IST/ULisboa) through grant No. UID/PRR/00099/2025 and grant No. UID/00099/2025. M.M.C. and C.F.B.M. acknowledge Fundação Amazônia de Amparo a Estudos e Pesquisas (FAPESPA), Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) – Finance Code 001, from Brazil, for partial financial support. R.P.M. acknowledges support from the Villum Investigator program supported by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. This project has received funding from the European Union’s Horizon MSCA-2022 research and innovation programme “Einstein Waves” under grant agreement No. 101131233.